OpenStudy (anonymous):

Find the surface area of a cone with a diameter of 13 and a slant height of 5. (use 3.14 for pi) 234.72 433.28 198.75 734.76

OpenStudy (anonymous):

The surface area of a cone is given by the formula: Sa=B+1/2*CL. Where B is the area of the base, C is the circumference , and L is the slant height.

OpenStudy (anonymous):

To find the area of the base, since this is a cone, we use the formula for the area of a circle: pi*r^2. In this case r is 6.5.

OpenStudy (anonymous):

The circumference is found using th formula: 2*pi*r. Again r=6.5

OpenStudy (anonymous):

The slant height is given as being 5

OpenStudy (anonymous):

so whats B?

OpenStudy (anonymous):

B is the area of the base, the base of a cone is a circle, so B=pi*6.5^2

OpenStudy (anonymous):

thank you :)